# Gravity from entropy

### v5

@@ Appendix B Eigenvalues and entropy for metrics ${\bf G}_{(1)}$ and for metrics ${\bf G}_{(n)}$
In this Appendix our goal is to use algebraic geometry [54] to define the eigenvalues of the metrics ${\bf G}_{(n)}$ considered as quantum opeartors between $n$-forms and their associated entropy. In order to establish the foundation of this treatment let us revisit the treatment outlined in Sec. II for defining the eigenvectors and the associated entropy of the metrics ${\bf G}_{(1)}$ viewed as qauntum operators between $1$-forms and treat on the same footing metrics tensor ${\bf G}_{(n)}$ viewed as quantum operators between $n$-forms with arbitrary values of $n$.
@@ B.1 Eigenvalues of metric matrices ${\bf G}_{(n)}$
${\bf G}_{(2)}\cdot\zeta=\tilde{\zeta}=\hat{\zeta}_{\mu\nu}dx^{\mu}\wedge dx^{\nu},$ (83)
@@
where here and in the following we have indicated with $\vec{\mu}$ $\vec{\rho}$ and $\vec{\eta}$ tuples of $n$ indices. From this definition of the eigenvector of a metric between $n$-forms it follows immediately that the metric tensor ${g}_{(n)}$ as all eigenvalues $\lambda$ equal to one, i.e.
@@ B.2 Flattening of the metric tensors ${\bf G}_{(n)}$ and reduction to a matrix eigenvalue problem
Similarly metric tensors ${\bf G}_{(n)}$ considered as quantum operators between $n$-forms can be flattened by considering a basis of independent $n$-forms. For instance a basis for these $n$ forms can be given by the canonical $n$-forms
@@
with $\mu_{1}<\mu_{2}<\ldots<\mu_{n}$. This canonical basis is formed by $m=\begin{pmatrix}d\\ n\end{pmatrix}$ forms.
@@ Appendix C The topological metric as a quantum operator
  • Scalar product of a topological field with itself The scalar product of a topological field with itself is real
@@
$+{[G_{(2)}]}_{\mu\nu\rho\sigma}\Big{(}dx^{\mu}\wedge dx^{\nu}\Big{)}\otimes\Big{(}dx^{\rho}\wedge dx^{\sigma}\Big{)}.$
@@
$+{[G_{(2)}]}_{\mu\nu\rho\sigma}\omega^{\rho}\omega^{\sigma}dx^{\mu}\wedge dx^{\nu}.$
@@
$+{[G_{(2)}^{\star}]}^{\mu\nu\rho\sigma}\Big{(}dx_{\mu}\wedge dx_{\nu}\Big{)}\otimes\Big{(}dx_{\rho}\wedge dx_{\sigma}\Big{)}.$
@@
$+{[G_{(2)}^{\star}]}^{\mu\nu\rho\sigma}\zeta_{\rho\sigma}dx_{\mu}\wedge dx_{\nu}.$
@@
${[G_{(2)}]}^{\mu\nu\rho\sigma}$ $=$ ${[g_{(2)}]}^{\mu\nu\eta_{1}\eta_{2}}{[G_{(2)}]}_{\eta_{1}\eta_{2}\theta_{1}\theta_{2}}{[g_{(2)}]}^{\rho\sigma\theta_{1}\theta_{2}}.$
@@
Interesting the metric tensors are such that the dual of the dual coincides with the metric tensor itself, indeed we have
@@
    $\hskip 11.38109pt+{[G_{(2),12}]}_{\mu\nu\rho\sigma}\Big{(}dx^{\mu}\wedge dx^{\nu}\Big{)}\otimes\Big{(}dx^{\rho}\wedge dx^{\sigma}\Big{)},$
@@
    $+{[G_{(2),12}^{\star}]}^{\mu\nu\rho\sigma}\Big{(}dx_{\mu}\wedge dx_{\nu}\Big{)}\otimes\Big{(}dx_{\rho}\wedge dx_{\sigma}\Big{)},$
@@
    $\mbox{Tr}_{F}\Big{(}\tilde{\bf G}\tilde{\bf G}^{\star}\Big{)}$ $=$ $G_{(0)}^{2}+{[G_{(1)}]}_{\mu\nu}{[G_{(1)}]}^{\mu\nu}$ (137)
    $+{[G_{(2)}]}_{\mu\nu\rho\sigma}{[G_{(2)}]}^{\mu\nu\rho\sigma}.$
@@
$+{[G_{(2)}]}_{\mu\nu\eta_{1}\eta_{2}}g_{(2)}^{\eta_{1}\eta_{2}\rho\sigma}\zeta_{\rho\sigma}\Big{(}dx^{\mu}\wedge dx^{\nu}\Big{)}.$

### v7

@@ Appendix B Eigenvalues and entropy for metrics ${\bf G}_{(1)}$ and for metrics ${\bf G}_{(n)}$
In this Appendix our goal is to use algebraic geometry [54] to define the eigenvalues of the metrics ${\bf G}_{(n)}$ considered as quantum operators between $n$-forms and their associated entropy. In order to establish the foundation of this treatment let us revisit the treatment outlined in Sec. II for defining the eigenvectors and the associated entropy of the metrics ${\bf G}_{(1)}$ viewed as quantum operators between $1$-forms and treat on the same footing metrics tensor ${\bf G}_{(n)}$ viewed as quantum operators between $n$-forms with arbitrary values of $n$.
@@ B.1 Eigenvalues of metric matrices ${\bf G}_{(n)}$
${\bf G}_{(2)}\cdot\zeta=\hat{\zeta}=\hat{\zeta}_{\mu\nu}dx^{\mu}\wedge dx^{\nu},$ (82)
@@
where here and in the following we have indicated with $\vec{\mu}$ $\vec{\rho}$ and $\vec{\eta}$ tuples of $n$ indices. From this definition of the eigenvector of a metric between $n$-forms it follows immediately that the metric tensor ${g}_{(n)}$ has all eigenvalues $\lambda$ equal to one, i.e.
@@ B.2 Flattening of the metric tensors ${\bf G}_{(n)}$ and reduction to a matrix eigenvalue problem
Similarly metric tensors ${\bf G}_{(n)}$ considered as quantum operators between $n$-forms can be flattened by considering a basis of independent $n$-forms. For instance a basis for these $n$-forms can be given by the canonical $n$-forms
@@
with $\mu_{1}<\mu_{2}<\ldots<\mu_{n}$. This canonical basis is formed by $m=\begin{pmatrix}d\\ n\end{pmatrix}$ $n$-forms.
@@ Appendix C The topological metric as a quantum operator
  • Scalar product of a topological field with itself The scalar product of a topological field with itself is real:
@@
$\oplus{[G_{(2)}]}_{\mu\nu\rho\sigma}\Big{(}dx^{\mu}\wedge dx^{\nu}\Big{)}\otimes\Big{(}dx^{\rho}\wedge dx^{\sigma}\Big{)}.$
@@
$\oplus{[G_{(2)}]}_{\mu\nu\rho\sigma}\zeta^{\rho\sigma}dx^{\mu}\wedge dx^{\nu}.$
@@
$\oplus{[G_{(2)}^{\star}]}^{\mu\nu\rho\sigma}\Big{(}dx_{\mu}\wedge dx_{\nu}\Big{)}\otimes\Big{(}dx_{\rho}\wedge dx_{\sigma}\Big{)}.$
@@
$\oplus{[G_{(2)}^{\star}]}^{\mu\nu\rho\sigma}\zeta_{\rho\sigma}dx_{\mu}\wedge dx_{\nu}.$
@@
${[G_{(2)}]}^{\mu\nu\rho\sigma}$ $=$ ${[g_{(2)}]}^{\mu\nu\eta_{1}\eta_{2}}{[G_{(2)}]}_{\eta_{1}\eta_{2}\theta_{1}\theta_{2}}{[g_{(2)}]}^{\theta_{1}\theta_{2}\rho\sigma}.$
@@
Interestingly, the metric tensors are such that the dual of the dual coincides with the metric tensor itself, indeed we have
@@
    $\hskip 11.38109pt\oplus{[G_{(2),12}]}_{\mu\nu\rho\sigma}\Big{(}dx^{\mu}\wedge dx^{\nu}\Big{)}\otimes\Big{(}dx^{\rho}\wedge dx^{\sigma}\Big{)},$
@@
    $\oplus{[G_{(2),12}^{\star}]}^{\mu\nu\rho\sigma}\Big{(}dx_{\mu}\wedge dx_{\nu}\Big{)}\otimes\Big{(}dx_{\rho}\wedge dx_{\sigma}\Big{)},$
@@
    $\mbox{Tr}_{F}\Big{(}\tilde{\bf G}\tilde{\bf G}^{\star}\Big{)}$ $=$ $G_{(0)}^{2}+{[G_{(1)}]}_{\mu\nu}{[G_{(1)}]}^{\nu\mu}$ (136)
    $+{[G_{(2)}]}_{\mu\nu\rho\sigma}{[G_{(2)}]}^{\rho\sigma\mu\nu}.$
@@
$\oplus{[G_{(2)}]}_{\mu\nu\eta_{1}\eta_{2}}g_{(2)}^{\eta_{1}\eta_{2}\rho\sigma}\zeta_{\rho\sigma}\Big{(}dx^{\mu}\wedge dx^{\nu}\Big{)}.$